OpenStudy (anonymous):

Solve. x^2 – 9x = –14 x = 2 and x = –7 x = 2 and x = 7 x = –2 and x = –7 x = –2 and x = 7

OpenStudy (mimi_x3):

(x-2)(x-7) x= 2,7

OpenStudy (anonymous):

x = 2 and x = 7

OpenStudy (anonymous):

x^2 – 9x = –14 x^2-9x+14=0 x^2-2x-7x+14=0 x(x-2)-7(x-2)=0 (x-2)(x-7)=0 =>x=2,7

OpenStudy (anonymous):

x = 2 and x = 7 x^2 – 9x = –14 x^2 – 9x +14=0 x^2-7x-2x+14=0 (x-7)(x-2)=0

OpenStudy (anonymous):

\[x ^{2}-9x=-14\]

OpenStudy (anonymous):

\[x ^{2}-9x+14=0\]

OpenStudy (anonymous):

(x-7)(x-2)=0 x=7,2